# Towards Graph Foundation Models: Learning Generalities Across Graphs via Task-Trees

### v2

@@ Abstract
Foundation models aim to create general, cross-task, and cross-domain machine learning models by pretraining on large-scale datasets to capture shared patterns or concepts, such as contours, colors, textures, and edges in images, or tokens, words, and sentences in text. However, identifying generalities across graph-structured data remains a significant challenge, as different graph-based tasks necessitate distinct inductive biases, thereby impeding the development of graph foundation models. To address this challenge, we introduce a novel approach for learning cross-task generalities in graphs. Specifically, we propose task-trees as basic learning instances to align task spaces (node, link, graph) on graphs. Then, we conduct a theoretical analysis to examine their stability, transferability, and generalization. Our findings indicate that when a graph neural network (GNN) is pretrained on diverse task-trees using a reconstruction objective, it acquires transferable knowledge, enabling effective adaptation to downstream tasks with an appropriate set of fine-tuning samples. To empirically validate this approach, we develop a pretrained graph model based on task-trees, termed Graph Generality Identifier on Task-Trees (GIT). Extensive experiments demonstrate that a single pretrained GIT model can be effectively adapted to over 30 different graphs across five domains via fine-tuning, in-context learning, or zero-shot learning. Our data and code are available at https://github.com/Zehong-Wang/GIT.
@@ 1 Introduction
Foundation models have gained prominence in the era of artificial general intelligence as a general-purpose, cross-task, and cross-domain machine learning approach. These models are exemplified by large language models (LLMs) for text (Achiam et al., 2023; Touvron et al., 2023) and large vision models (LVMs) for images (He et al., 2022; Yuan et al., 2021). Pretrained on large-scale datasets, they capture generalizable and transferable knowledge, including contours, colors, textures, and edges in images, as well as tokens, words, and sentences in text. These patterns represent modality-specific generalities, allowing the models to adapt to new downstream tasks via in-context learning (Xie et al., 2022) or zero-shot learning (Wei et al., 2021).
Despite the success of foundation models across various modalities, their development of graph-structured data remains in its infancy (Liu et al., 2024). This is primarily due to the high degree of variability among graph-structured datasets (Mao et al., 2024), making it challenging to identify shared generalities across graphs. In particular, graphs from different domains typically represent distinct phenomena; for example, social networks capture relationships between people (Freeman, 2004), while molecular networks depict the structures of molecules (Zeng et al., 2022). These distinctions are evident not only in feature heterogeneity (Wang et al., 2024b) but also in structural heterogeneity (Qiu et al., 2020; Wang et al., 2024c). More importantly, different graph-related tasks often do not share basic learning instances—such as nodes in node-level tasks and entire graphs in graph-level tasks—making it difficult to use a single model for diverse tasks (Wang et al., 2024b). The above challenges make it difficult to develop a foundation model that can identify generalities applicable across graphs. We focus on the last problem, i.e., task heterogeneity.
Is it possible to identify cross-task generalities across graphs? Despite the challenges, some researchers have attempted to uncover shared generalities across graphs, which can be mainly categorized into two groups. (1) One line of works draws on graph theory. In particular, they use the concept of graphon (Ruiz et al., 2020) to describe transferable patterns across graphs. If two graphs are generated from the same graphon, they are expected to share similar topological properties, which can result in high transferability between them (Ruiz et al., 2020; Cao et al., 2023). However, the strong assumptions underlying graphon theory are often unrealistic for real-world graphs (Levie et al., 2021), and even when the assumptions hold, finding the graphon shared from a large set of graphs presents another significant challenge. (2) Another approach involves leveraging substructures that are present across graphs, such as triangles, stars, and $k$-cliques (Zhao et al., 2023; Mao et al., 2024). For instance, triangle structures frequently appear in citation networks, social networks, and molecular networks, albeit with varying semantics. Based on this observation, some works (Sun et al., 2023; Liu et al., 2024) sample subgraphs consisting of substructures and then use a graph neural network (GNN) to encode the subgraph-level embeddings for prediction. However, message-passing GNNs are proved to struggle with capturing such basic substructures in learning (sub)graph embeddings (Garg et al., 2020; Esser et al., 2021; Zhang et al., 2024a), limiting the practicality of these methods.
@@
Given the limitations of existing approaches, we propose a novel perspective to answer the question by focusing on the learning process of GNNs. In message-passing GNNs (Kipf & Welling, 2017; Hamilton et al., 2017), predictions are made based on the so-called task-relevant nodes in the graph. For node-level tasks, the task-relevant node is the node itself; for edge-level tasks, it includes the start and end nodes of the edge; and for graph-level tasks, all nodes in the graph are task-relevant. Regardless of the task, basic GNN models only aggregate the embeddings of task-relevant nodes for predictions (Srinivasan & Ribeiro, 2020). This aggregation process is analogous to introducing a virtual “task node” that connects all task-relevant nodes, and learn the embedding of the task nodes for prediction. We refer the computation tree surrounding the “task node” as task-tree, as illustrated in Figure 1. The task-tree formulation offers three distinct advantages. (1) Learnability: the information within tree structures can be fully captured by message-passing GNNs (Gupta et al., 2024). (2) Uniformity: task-trees are applicable to node-, edge-, and graph-level tasks, preventing the task heterogeneity on graphs. (3) Efficiency: encoding task-trees technically equals to learning the embeddings of virtual task nodes appended to the original graph, without the need of extracting subgraphs. Building on these benefits, task-trees are able to serve as the basic learning instances in graphs, similar to images in vision and sentences in language. Therefore, task-trees may maintain transferable patterns or concepts shared across different graph tasks, leading to the following assumption:
@@
Task-Tree Generality Assumption: the generalities shared across graphs are (at least partially) preserved within the task-trees of the involved graphs.
@@
To validate this assumption, we conduct a theoretical analysis to examine the properties of task-trees in terms of stability, transferability, and generalization. Our key finding is that if a GNN model is pretrained on diverse task-trees using a reconstruction task, it will acquire sufficient transferable knowledge to downstream tasks with a reasonable number of fine-tuning samples. Additionally, the analysis highlights the potential for specializing the pretrained model into specific domains through post-training (Wei et al., 2021) using domain-specific instances.
@@
To empirically verify our theoretical findings, we propose Graph Generality Identifier on Task-Trees (GIT), which is pretrained on task-trees extracted from a wide range of graphs spanning diverse domains and tasks. We evaluate GIT across 32 graphs from 5 distinct domains, via fine-tuning, in-context learning (few-shot learning without fine-tuning), and zero-shot learning. The experimental results demonstrate that pretraining the model on a small set of graphs significantly benefits a wide range of graph tasks, supporting the idea that task-trees indeed capture generalities shared across graph-based tasks. We also propose an instruction tuning method to adapt the pretrained general model to specific downstream domains, which enhances performance across all datasets within the domain, matching or even surpassing domain expert models. We summarize our contributions as follows:
@@
  • We introduce task-trees as basic learning instances for aligning diverse graph-based tasks, demonstrating their superiority over existing instances such as subgraphs.
@@
  • We provide a comprehensive theoretical analysis of the effectiveness of task-trees as the learning instances for task alignment. To the best of our knowledge, this is the first theoretical framework addressing task heterogeneity in graph learning.
  • Leveraging task-trees, we propose GIT, a pretrained model that captures generalizable knowledge across graphs and can be adapted to specialized domains.
@@
  • Extensive experiments conducted on 32 diverse graphs validate the effectiveness of GIT across multiple learning paradigms, including fine-tuning, in-context learning, and zero-shot learning.
2 Task-Trees: Rethinking Basic Learning Tokens in Graph Task Alignment
@@ 2.2 Task-Tree Construction and Encoding
The learning process of GNNs involves recursively integrating information from the computation trees, progressing from the bottom to the top. Therefore, for a given node $v$, the node embedding produced by an $L$-layer GNN corresponds to the embedding of its computation tree $T_{v}^{L}$. Since the prediction for any graph task depends solely on the embeddings of task-relevant nodes, and these embeddings are determined by their respective computation trees, we can construct a task-tree for each instance—whether it be a node, edge, or graph—as illustrated in Figure 1.
@@
We adopt a straightforward approach to encode task-trees. Given a task-tree $T^{t}$, which consists of a virtual task node $v^{t}$ and a set of task-relevant nodes $\{v_{1}^{t},\dots,v_{n}^{t}\}$, we employ a MEAN aggregator for encoding:
@@
where $T_{i}$ represents the computation tree associated with the task-relevant node $v_{i}^{t}$, and $\phi$ denotes the GNN encoder.
@@
2.3 Comparison to Existing Works
Unlike our proposed task-trees, several existing approaches (Qiu et al., 2020; Sun et al., 2023; Huang et al., 2023; Liu et al., 2024; He & Hooi, 2024) utilize $k$-hop subgraphs extracted from graphs as the basic learning instances. For instance, in node classification, ego-graphs are constructed around each node, where the label of the central node is assigned to the induced subgraph, effectively reformulating node classification as a subgraph classification task. A similar transformation can be applied to edge-level and graph-level tasks by converting them into subgraph-level learning problems. This method involves: (1) extracting ego-graphs centered around task-relevant nodes and (2) applying GNNs to learn graph-level embeddings for classification. However, this subgraph extraction process incurs substantial computational overhead, increasing both time and memory requirements due to the necessity of storing and processing induced subgraphs. Moreover, information within these subgraphs is not always effectively captured by message-passing GNNs, as GNNs may struggle to learn essential substructures preserved in graphs (Garg et al., 2020; Chen et al., 2020; Zhang et al., 2024a), thereby limiting the efficacy of subgraphs as learning instances.
In contrast, our proposed task-trees offer both greater efficiency and improved learnability. Specifically, encoding task-trees for node, link, or graph-level tasks involves: (1) augmenting the original graph by adding virtual task nodes and connecting them to task-relevant nodes and (2) applying GNNs over the augmented graph to encode the embeddings of these virtual nodes for prediction. For example, in node classification, we first introduce virtual nodes connected to each node in the original graph and subsequently apply GNNs to this augmented structure, allowing virtual node embeddings to be learned for classification. Consequently, our method requires only the addition of nodes and edges to the existing graph, making it significantly more efficient than extracting and storing subgraphs. Furthermore, encoding task-trees is equivalent to directly encoding virtual nodes through message passing, ensuring that task-tree information remains fully learnable by standard GNNs. Empirically, task-trees consistently outperform subgraphs in both effectiveness and efficiency (Section 5.5 and Appendix I). A detailed comparison is presented in Table 1, with additional discussions on related work available in Appendix A.
3 Why Do Task-Trees Serve as Basic Learning Instances?
In this section, we provide a theoretical analysis of task-trees, highlighting their stability, transferability, and generalization properties as basic learning instances. This analysis serves to validate the Task-Tree Generality Assumption from a theoretical standpoint. Importantly, our objective is not to claim the superiority of task-trees over other approaches, such as subgraphs. Rather, we establish the feasibility of learning cross-task generalities based on task-trees.
@@ 3 Why Do Task-Trees Serve as Basic Learning Instances?
Theorem 3.1 (proved in Appendix D.1) suggests that two task-trees are likely to have similar representations if their subtrees are similar. This theorem highlights the significance of similarity between pairs of subtrees, while downplaying the impact of the number of subtrees (i.e., the width of the task-trees), despite having more subtrees could potentially increase diversity and thus magnify discrepancy. The theorem also implies that increasing the number of GNN layers may lead to a loose bound, which aligns with previous analyses (Garg et al., 2020; Ju et al., 2023).
@@ 4 Graph Generality Identifier on Task-Trees
The theoretical analysis demonstrates the feasibility of building graph foundation models based on task-trees. We apply these theorems to develop our GIT model, with the aim of empirically validating the Task-Tree Generality Assumption.
As our focus is to align task spaces on graphs, we adopt existing text-attributed graph benchmarks (Chen et al., 2024b; Zhang et al., 2024b; Feng et al., 2024) to enable feature space alignment. In particular, we use Sentence Bert (Reimers & Gurevych, 2019) to convert node features of all graphs into 768-dim vectors, following Liu et al. (2024).
@@ 4.1 GIT-G: Pretraining to Acquire General Knowledge
We propose a task-tree reconstruction task as a pretext for pretraining. The key is to use two corrupted task-trees to reconstruct each other, thereby capturing corruption-invariant semantics. Given a set of task-trees $\{T^{t}_{1},...,T^{t}_{n}\}$ sampled from a graph database, we apply corruption techniques to generate two views of each task-tree, denoted as $\{\hat{T}^{t}_{1},...,\hat{T}^{t}_{n}\}$ and $\{\tilde{T}^{t}_{1},...,\tilde{T}^{t}_{n}\}$. For corruption, we use random edge masking and random attribute masking (Zhu et al., 2020) due to its computational efficiency. We then use an encoder $\phi$ to obtain embeddings for the corrupted task-trees, resulting in $\{\hat{{\bm{z}}}_{1},...,\hat{{\bm{z}}}_{n}\}$ and $\{\tilde{{\bm{z}}}_{1},...,\tilde{{\bm{z}}}_{n}\}$. Inspired by (Thakoor et al., 2022), we perform reconstruction as

### v3

@@ Abstract
Foundation models are pretrained on large-scale corpora to learn generalizable patterns across domains and tasks—such as contours, textures, and edges in images, or tokens and sentences in text. In contrast, discovering such generalities in graph-structured data, especially across heterogeneous graph tasks, remains an open challenge. To address this, we propose a novel approach to cross-task generalization in graphs via task-trees, which serve as unified learning instances aligning node-, edge-, and graph-level tasks. We theoretically analyze the stability, transferability, and generalization properties of task-trees, showing that pretraining a graph neural network (GNN) on diverse task-trees with a reconstruction objective induces transferable knowledge. This enables efficient adaptation to downstream tasks with minimal fine-tuning. To validate our framework, we introduce Graph Generality Identifier on Task-Trees (GIT), a graph foundation model that demonstrates strong performance on over 30 graphs across five domains via fine-tuning, in-context learning, and zero-shot generalization. Code and data are available at https://github.com/Zehong-Wang/GIT.
Machine Learning, ICML
@@ 1 Introduction
Foundation models have emerged as a cornerstone of general-purpose machine learning, enabling cross-task and cross-domain generalization. Representative examples include large language models (LLMs) for text (Achiam et al., 2023; Touvron et al., 2023) and large vision models (LVMs) for images (He et al., 2022; Yuan et al., 2021). Pretrained on massive datasets, these models capture transferable patterns—such as contours and textures in images, or tokens and sentences in text—that reflect modality-specific generalities. This broad knowledge base allows for efficient adaptation to downstream tasks via in-context learning (Xie et al., 2022; Chen et al., 2024c) and zero-shot generalization (Wei et al., 2021).
Despite the success of foundation models in text and vision, their extension to graph-structured data remains nascent (Liu et al., 2024a), primarily due to the high variability across graph datasets (Mao et al., 2024). Graphs from different domains often encode distinct phenomena—e.g., social networks model human relationships (Freeman, 2004), whereas molecular graphs represent chemical structures (Zeng et al., 2022)—leading to both feature (Wang et al., 2024b) and structural heterogeneity (Qiu et al., 2020; Wang et al., 2024c). Crucially, graph tasks operate on different learning units, such as nodes, edges, or entire graphs, limiting cross-task compatibility within a unified model (Wang et al., 2024b). These challenges hinder the development of graph foundation models capable of capturing transferable generalities. In this work, we specifically address the challenge of task heterogeneity.
Is it possible to identify cross-task generalities across graphs? Despite inherent challenges, prior work has explored this question via two main approaches. (1) A graph-theoretic perspective employs the concept of graphons (Ruiz et al., 2020) to model transferable patterns across graphs. If graphs are sampled from the same graphon, they are expected to share structural properties, enabling effective transfer (Ruiz et al., 2020; Cao et al., 2023). However, graphon-based methods rely on strong generative assumptions that rarely hold in real-world settings (Levie et al., 2021), and inferring a shared graphon from diverse graphs remains computationally intractable. (2) A substructure-based perspective seeks recurring motifs—such as triangles, stars, or $k$-cliques—across domains (Zhao et al., 2023; Mao et al., 2024). These motifs appear in various contexts (e.g., social, citation, and molecular networks), motivating methods that sample subgraphs consisting of substructures and encode them via GNNs (Sun et al., 2023; Liu et al., 2024a). However, message-passing GNNs are fundamentally limited in capturing such substructures (Garg et al., 2020; Esser et al., 2021; Zhang et al., 2024a), restricting their efficacy in learning transferable subgraph representations.
@@
Given the limitations of prior approaches, we introduce a novel perspective centered on the learning dynamics of message-passing GNNs (Kipf & Welling, 2017; Hamilton et al., 2017). In such models, predictions are made based on task-relevant nodes: the target node in node-level tasks, edge endpoints in edge-level tasks, and all nodes in graph-level tasks (Srinivasan & Ribeiro, 2020). Regardless of task type, GNNs aggregate embeddings over these task-relevant nodes, which can be conceptualized as introducing a virtual task node connected to all task-relevant nodes. We define the computation tree rooted at this virtual node as a task-tree (Figure 1). Task-trees offer three key advantages: (1) Learnability: tree-structured information can be effectively captured by message-passing GNNs (Gupta et al., 2024); (2) Uniformity: task-trees apply seamlessly across node-, edge-, and graph-level tasks, mitigating task heterogeneity; (3) Efficiency: encoding task-trees operationally equals to encoding the virtual nodes appended to original graphs. Analogous to images in vision or sentences in language, task-trees serve as unified learning instances and may encode transferable patterns across graph tasks, leading to the assumption:
@@
Task-Tree Generality Assumption. The generalities shared across graphs are (at least partially) preserved within the task-trees of the involved graphs.
@@
To evaluate this assumption, we conduct a theoretical analysis of task-trees with respect to stability, transferability, and generalization. Our main result shows that pretraining a GNN on diverse task-trees via a reconstruction objective yields transferable representations that adapt well to downstream tasks with moderate fine-tuning. Furthermore, the model can be specialized to specific domains via post-training (Wei et al., 2021) on domain-specific task-trees.
@@
To empirically validate our theoretical insights, we introduce Graph Generality Identifier on Task-Trees (GIT), a graph foundation model pretrained on task-trees extracted from diverse graphs spanning multiple domains and tasks. GIT is evaluated on 32 graphs across 5 domains under three paradigms: fine-tuning, in-context learning (few-shot without fine-tuning), and zero-shot learning. Results show that pretraining on a small set of graphs significantly improves performance on a broad range of downstream tasks, supporting the hypothesis that task-trees capture transferable generalities. Additionally, we propose an instruction tuning method to adapt the general model to specific domains, yielding performance comparable to or exceeding domain-specific expert models. Our key contributions are:
@@
  • We introduce task-trees as unified learning instances for aligning heterogeneous graph tasks, demonstrating advantages over conventional units such as subgraphs.
@@
  • We present the first theoretical framework addressing task heterogeneity in graph learning, establishing the effectiveness of task-trees for cross-task generalization.
  • We propose GIT, a graph foundation model pretrained on task-trees to acquire generalizable knowledge and support domain specialization.
@@
  • Extensive experiments across 32 graphs and five domains validate the effectiveness of GIT under fine-tuning, in-context, and zero-shot settings.
2 Task-Trees: Rethinking Basic Learning Instances on Graphs
@@ 2.2 Task-Tree Construction and Encoding
The learning process of message-passing GNNs can be interpreted as recursive aggregation over computation trees, where the representation of a node $v$ produced by an $L$-layer GNN corresponds to the embedding of its $L$-hop computation tree $T_{v}^{L}$. Since predictions in graph tasks rely exclusively on the embeddings of task-relevant nodes, and those embeddings are determined by their respective computation trees, we construct a unified task-tree for each learning instance—whether a node, edge, or entire graph—by merging the relevant computation trees, as illustrated in Figure 1.
@@
To encode a task-tree, we adopt a simple yet effective aggregation strategy. Given a task-tree $T^{t}$ composed of a virtual node $v^{t}$ and task-relevant nodes $\{v_{1}^{t},\dots,v_{n}^{t}\}$, we compute its representation using a MEAN aggregator over the embeddings of the individual computation trees:
@@
where $T_{i}$ denotes the computation tree rooted at $v_{i}^{t}$, and $\phi$ is a shared GNN encoder. This representation serves as the input for downstream objectives such as reconstruction, classification, or alignment.
2.3 Comparison to Existing Works
Unlike our proposed task-trees, several existing approaches (Qiu et al., 2020; Sun et al., 2023; Huang et al., 2023; Liu et al., 2024a; He & Hooi, 2024) utilize $k$-hop subgraphs extracted from graphs as the basic learning instances. For instance, in node classification, ego-graphs are constructed around each node, where the label of the central node is assigned to the induced subgraph, effectively reformulating node classification as a subgraph classification task. A similar transformation can be applied to edge-level and graph-level tasks by converting them into subgraph-level learning problems. This method involves: (1) extracting ego-graphs centered around task-relevant nodes and (2) applying GNNs to learn graph-level embeddings for classification. However, this subgraph extraction process incurs substantial computational overhead, increasing both time and memory requirements due to the necessity of storing and processing induced subgraphs. Moreover, information within these subgraphs is not always effectively captured by message-passing GNNs, as GNNs may struggle to learn essential substructures preserved in graphs (Garg et al., 2020; Chen et al., 2020; Zhang et al., 2024a), thereby limiting the efficacy of subgraphs as learning instances.
In contrast, our proposed task-trees offer both greater efficiency and improved learnability (Table 1). Specifically, encoding task-trees for node, link, or graph-level tasks involves: (1) augmenting the original graph by adding virtual task nodes and connecting them to task-relevant nodes and (2) applying GNNs over the augmented graph to encode the embeddings of these virtual nodes for prediction. For example, in node classification, we first introduce virtual nodes connected to each node in the original graph and subsequently apply GNNs to this augmented structure, allowing virtual node embeddings to be learned for classification. Consequently, our method requires only the addition of nodes and edges to the existing graph, making it significantly more efficient than extracting and storing subgraphs. Furthermore, encoding task-trees is equivalent to directly encoding virtual nodes through message passing, ensuring that task-tree information remains fully learnable by standard GNNs. Empirically, task-trees consistently outperform subgraphs in both effectiveness and efficiency (Section 5.6 and Appendix I).
@@ 2.3 Comparison to Existing Works
The most closely related work to our task-tree framework is GFT (Wang et al., 2024b), which introduces computation trees to align heterogeneous graph tasks. While both approaches share the core intuition of structuring task-specific trees, GFT adopts a model-centric perspective, featuring a learnable vocabulary, a multi-faceted reconstruction objective, and specialized adaptation classifiers. In contrast, our task-tree framework is theory-driven and emphasizes the design of learning instances rather than model complexity. Notably, GFT empirically demonstrates the potential of computation trees for transferability, while our work complements it with a formal theoretical foundation, offering a principled understanding of task alignment in graph learning. Together, these works represent complementary advances toward general-purpose graph foundation models. We present additional discussions on related work in Appendix A.
@@
3 Theoretical Analysis of Task-Trees
In this section, we present a theoretical analysis of task-trees, focusing on their stability, transferability, and generalization as foundational learning instances. This analysis provides formal support for the Task-Tree Generality Assumption, which posits that transferable patterns across graph tasks are preserved within task-tree structures.
Our goal is not to assert the universal superiority of task-trees over other learning units such as subgraphs, but rather to establish the theoretical plausibility of using task-trees to capture cross-task generalities. By grounding the construction and use of task-trees in formal guarantees, we lay the foundation for principled pretraining and transfer learning across heterogeneous graph tasks.
@@ 3 Theoretical Analysis of Task-Trees
Theorem 3.1 (proved in Appendix D.1) suggests that two task-trees are likely to have similar representations if their subtrees are similar. This theorem highlights the significance of similarity between pairs of subtrees, while downplaying the impact of the number of subtrees (i.e., the width of the task-trees), despite having more subtrees could potentially increase diversity and thus magnify discrepancy. The theorem also implies that increasing the number of GNN layers may lead to a loose bound, which aligns with previous analyses (Garg et al., 2020; Ju et al., 2023a).
@@ 4 Graph Generality Identifier on Task-Trees
The theoretical analysis establishes the feasibility of constructing graph foundation models based on task-trees. Building on these insights, we develop the GIT model to empirically validate the Task-Tree Generality Assumption.
To focus on aligning task spaces across heterogeneous graph tasks, we adopt widely used text-attributed graph benchmarks (Chen et al., 2024b; Zhang et al., 2024b; Feng et al., 2024), which simplify feature alignment across datasets. Specifically, we follow Liu et al. (2024a) and encode all node features into a shared 768-dimensional embedding space using Sentence-BERT (Reimers & Gurevych, 2019). This design allows us to isolate and examine the effect of task-tree-based pretraining while holding node features consistent across domains.
@@ 4.1 GIT-G: Pretraining to Acquire General Knowledge
We propose a task-tree reconstruction task as a pretext for pretraining. The key is to use two corrupted task-trees to reconstruct each other, thereby capturing corruption-invariant semantics. Given a set of task-trees $\{T^{t}_{1},...,T^{t}_{n}\}$ sampled from a graph database, we apply corruption techniques to generate two views of each task-tree, denoted as $\{\hat{T}^{t}_{1},...,\hat{T}^{t}_{n}\}$ and $\{\tilde{T}^{t}_{1},...,\tilde{T}^{t}_{n}\}$. For corruption, we use random edge masking and random attribute masking (Zhu et al., 2020; Wang et al., 2025c) due to its computational efficiency. We then use an encoder $\phi$ to obtain embeddings for the corrupted task-trees, resulting in $\{\hat{{\bm{z}}}_{1},...,\hat{{\bm{z}}}_{n}\}$ and $\{\tilde{{\bm{z}}}_{1},...,\tilde{{\bm{z}}}_{n}\}$. Inspired by (Thakoor et al., 2022), we perform reconstruction as